OpenStudy (mtalhahassan2):

Need help!! 1. Determine the angular velocity of the hour hand of a clock in radians per second. (two significant figures please).

OpenStudy (mtalhahassan2):

@Mertsj

OpenStudy (mertsj):

\[\frac{2\pi}{12 hr}\times \frac{1 hr}{3600 \sec}=1.5 radians/\sec\]

OpenStudy (mtalhahassan2):

@Mertsj can you plz explain me how you get that values plz

OpenStudy (mertsj):

Let me check. Maybe I hit a wrong button.

OpenStudy (mtalhahassan2):

ok

OpenStudy (mertsj):

Whoops. 1.5 x10^-4

OpenStudy (mtalhahassan2):

oh ok but can you plz explain how you get 2pie/ 12 hrs and etc.

OpenStudy (mtalhahassan2):

I am confuse

OpenStudy (mertsj):

If the hour hand is on 1 where will it be in 12 hours?

OpenStudy (mtalhahassan2):

on 1 again

OpenStudy (mtalhahassan2):

because 12+1 is 13

OpenStudy (mertsj):

Yes. It will have moved all the way around the circle (2pi radians) back to its starting point in 12 hours.The rest is just unit conversion.

OpenStudy (mtalhahassan2):

ok but I still don't get 3600 how you get that

OpenStudy (mertsj):

How many seconds are in 1 hour where you live?

OpenStudy (mtalhahassan2):

60 sec

OpenStudy (mertsj):

In America there are 60 seconds in 1 minute

OpenStudy (mtalhahassan2):

so it be in 1 hour there is like 60min and 1 min is 60 sec. So it be 60X60 which is 3600

OpenStudy (mtalhahassan2):

right?

OpenStudy (mertsj):

Bingo!!! Good for you!!!

OpenStudy (mtalhahassan2):

thanks

OpenStudy (mertsj):

yw